# Higher-Order Regularization Learning on Hypergraphs

### v1

@@ 3.4.1 Fully supervised problem with HOHL regularization
$\|u_{n,\tau}^{(\mathbf{y}_{n})}-g|_{\Omega_{n}}\|_{\mathrm{L}^{2}(\mu_{n})}$ $\leq C\Bigg[\sum_{k=1}^{q}\lambda_{k}\frac{\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\left(\varepsilon_{n}^{(k)}\right)^{2p_{k}}}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}+\frac{\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}+\tau\left(1+\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}\right)$
$+\left(\frac{\log(n)}{n}\right)^{1/d}\Bigg]$
where $u_{n,\tau}^{(\mathbf{g}_{n})}$ is the minimizer of $\mathcal{R}_{n,\tau}^{(\mathbf{y}_{n})}$.
This result highlights several important aspects of the behavior of the discrete minimizers. First, the convergence rate explicitly depends on the interplay between the length-scales $\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ and the regularization parameter $\tau$, reflecting the multiscale nature of the HOHL regularizer. This dependence can guide practitioners in the choice of these parameters to balance bias, variance, and computational cost in practical applications. Second, the theorem generalizes previously known rates for graph-based learning: when $q=1$, we recover the convergence rates established in [34, Corollary 1.8], thus placing our result within and extending the existing theoretical framework on graphs.
@@ 3.4.2 Truncated energies
Next, we show that we can use the truncated version of HOHL in practice. In fact, going beyond heuristics, the below results shows that truncated energies converges to the same continuum energy as the full energy (see [85, Theorem 3.5]) This signifies that for large enough $n$, truncated and full energies will lead to arbitrarily close minimizers.
@@
Moreover, the continuum energy $\sum_{i=1}^{\infty}\Bigl(\sum_{r=1}^{q}\beta_{i}^{p_{r}}\Bigr)\langle v,\psi_{i}\rangle_{\mathrm{L}^{2}(\mu)}^{2}$ coincides with $\langle v,\mathcal{L}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu)}$, where $\mathcal{L}^{(q)}$ is an operator derived from $\Delta_{\rho}$. In particular, $\mathcal{L}^{(q)}$ is defined spectrally via its eigenfunctions and eigenvalues (see Lemma 4.6): it shares the same eigenfunctions as $\Delta_{\rho}$, while its eigenvalues are given by functions of those of $\Delta_{\rho}$. This places our approach firmly within the framework of spectral kernel learning, where it is common to regularize with operators derived from the Laplacian. Spectral learning—and its analysis through reproducing kernel Hilbert space techniques—has been shown to yield powerful results for uncertainty quantification, enabling explicit bounds on expected error as well as estimates of prediction variance in semi-supervised learning (see [91] and references therein).
@@ 3.4.3 Non-geometrical setting
All of the preceding results focused on applying HOHL within the geometric setting. The following result extends the analysis to arbitrary hypergraphs, demonstrating that the matrix $\mathcal{L}^{(q)}_{\mathrm{dis}}$ can be interpreted as the Laplacian of a specially constructed graph (which may be signed [76]).
@@
This result is particularly noteworthy as it implies that standard numerical techniques developed for Laplace learning are directly applicable to HOHL. These include spectral truncation (see Theorem 3.2), Nyström extensions [28], conjugate gradient methods for Laplacian inversion, and more. Moreover, it suggests that HOHL can function as a drop-in replacement for Laplace learning within existing machine learning pipelines. To demonstrate this in practice, Section 5 presents active learning experiments where the HOHL matrix $\mathcal{L}^{(q)}_{n}$ defines a Gaussian prior over functions.
We can extend the HOHL energy (2) to non-geometric datasets, where geometric embeddings for the vertices are unavailable and, for example, the weight models described in Section 3.2 do not apply. In such settings, the standard feature-based hypergraph construction, e.g. [93, 44], forms a hyperedge among all nodes that share a common categorical feature value. Each hyperedge is also assigned unit weight.
@@
Continuum limit of HOHL $\mathrm{W}^{p_{q},2}$ semi-norm [85] No natural continuum limit
Characterization of well/ill-posedness of HOHL in SSL ✓([85, Theorem 3.5])
@@ 4.1 Fully supervised problem with HOHL regularization
First, the aim is to show the analogue of [34, Proposition 2.1] and to this purpose, we define
@@
Let $E=\cap_{r=1}^{q}E_{r}$ be the set of events such that the above inequalities hold for all $\leq r\leq q$. By [34, Lemma 2.2], we know that $\mathbb{P}(E_{r})\geq 1-2ne^{-c(r)n\left(\varepsilon_{n}^{(r)}\right)^{d}}$. Hence,
@@
where we used the fact that $\left(W_{n,\varepsilon_{n}^{(k)}}\right)_{ij}\leq C\left(\varepsilon_{n}^{(k)}\right)^{-d}$ and (7) for (9) and the fact that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ for (10). This implies that $n^{\ell-1}(\varepsilon_{n}^{(1)})]^{-1}q_{j}^{i}$ are sub-Gaussian and satisfy the same inequalities in the Birnbaum-Orlicz norm as in [34, Lemma 2.4]. By applying the same Hoeffding inequality as in the latter, for any $t>0$, we therefore obtain
@@
The term $T_{-r_{\chi}}$ contains $p_{k}-r_{\chi}$ factors, each equal to either $D_{n,\varepsilon_{n}^{(k)}}$ or $W_{n,\varepsilon_{n}^{(k)}}$. Using the operator-norm bounds from [34, Lemma 2.3], we have $\|T_{-r_{\chi}}\|_{\mathrm{op}}\leq(Cn)^{p_{k}-r_{\chi}}$. This implies that
@@
$=\left\|\left[\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\prod_{i=1}^{p_{k}}D_{n,\varepsilon_{n}^{(k)}}^{\chi_{i}}(-W_{n,\varepsilon_{n}^{(k)}})^{1-\chi_{i}}\right]\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
$\leq\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\left\|\prod_{i=1}^{p_{k}}D_{n,\varepsilon_{n}^{(k)}}^{\chi_{i}}(-W_{n,\varepsilon_{n}^{(k)}})^{1-\chi_{i}}\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
$=\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\left\|\prod_{i=1}^{p_{k}}D_{n,\varepsilon_{n}^{(k)}}^{\chi_{i}}(-W_{n,\varepsilon_{n}^{(k)}})^{1-\chi_{i}}\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
$\leq\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\|T_{-r_{\chi}}\|_{\mathrm{op}}\left\|W_{n,\varepsilon_{n}^{(k)}}D_{n,\varepsilon_{n}^{(k)}}^{r_{\chi}-1}\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
(19) $\leq Cn^{p_{k}-r_{\chi}}\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\frac{n^{r_{\chi}}\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}$
(20) $=C(2^{p_{k}}-1)\frac{n^{p_{k}}\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}$
@@
where we used the fact that there exists $C_{1}\leq n^{-1}d_{n,i,\varepsilon_{n}^{(r)}}\leq C_{2}$ for all $1\leq i\leq n$ and $1\leq r\leq q$ for (22) and the fact that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ for (23). Let $A$ be the event such that (14) holds. Then, arguing as in [35, Lemma 2.6], we can show that $\mathbb{P}(A\,|\,E)\geq 1-Cn^{-\alpha}$. Analogously to the proof of Lemma 4.1, we conclude that $\mathbb{P}(A)\geq 1-Cn^{-\alpha}$. ∎
@@
$\langle\frac{1}{2}\nabla\mathcal{R}_{n,\tau}^{(\mathbf{a}_{n})}(v^{(1)}_{n})-\frac{1}{2}\nabla\mathcal{R}_{n,\tau}^{(\mathbf{a}_{n})}(v_{n}^{(2)}),v_{n}^{(1)}-v_{n}^{(2)}$ $\rangle_{\mathrm{L}^{2}(\mu_{n})}=\|v_{n}^{(1)}-v_{n}^{(2)}\|_{\mathrm{L}^{2}(\mu_{n})}^{2}$
$+\tau\sum_{k=1}^{q}\lambda_{k}\left\langle\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}\left(v_{n}^{(1)}-v_{n}^{(2)}\right),v_{n}^{(1)}-v_{n}^{(2)}\right\rangle_{\mathrm{L}^{2}(\mu_{n})}.$
@@
Furthermore, by first order optimality condition and (15), we have that
@@
where $u_{n,\tau}^{(\mathbf{g}_{n})}$ and $u_{\tau}$ are the minimizers of $\mathcal{R}_{n,\tau}^{(\mathbf{y}_{n})}$ and $\mathcal{R}_{\infty,\tau}^{(g)}$ respectively.
@@
then $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq\|\mathbf{a}_{n}\|_{\mathrm{L}^{2}(\mu_{n})}$. Indeed, by Proposition 3.3, we know that $\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}$ is a graph Laplacian, so we can apply the same proof as in [34, Lemma 2.14] with the eigenpairs of $\mathcal{L}_{n}^{(q)}=\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}$ to deduce (30).
@@
Let $E_{k}$ be the event such that [34, Theorem 2.8] holds for $\varepsilon_{n}^{(k)}$: we have
@@
where we used the fact that $u_{\tau}|_{\Omega_{n}}-u_{n,\tau}^{(\mathbf{g}_{n})}$ satisfies (30) with $\mathbf{a}_{n}=\tau\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}-\sum_{k=1}^{q}\lambda_{k}\Delta_{\rho}^{p_{k}}\right)u_{\tau}|_{\Omega_{n}}$ for (35) and [34, Theorem 2.8] for (36).
@@
where we used (38) for (39). By [22, Lemma 17], there exists $c$ and $C$ such that $c\|h\|_{\mathrm{W}^{s,2}(\Omega)}\leq\|h\|_{\mathcal{H}^{s}(\Omega)}\leq C\|h\|_{\mathrm{W}^{s,2}(\Omega)}$ for all $h\in\mathcal{H}^{s}(\Omega)$. From the above, we therefore deduce that $\|u_{\tau}\|_{\mathrm{W}^{s,2}(\Omega)}\leq C\|g\|_{\mathrm{W}^{s,2}(\Omega)}$. Finally, by Morrey’s inequality [53], for $s$ sufficiently large there exists C>0C’>0 such that
@@
(42) $=\langle\nabla\mathcal{R}_{\infty}^{(g)}(g),g-u_{\tau}\rangle_{\mathrm{L}^{2}(\mu)}-\|g-u_{\tau}\|_{\mathrm{L}^{2}(\mu)}^{2}-\tau\sum_{k=1}^{q}\lambda_{k}\langle g-u_{\tau},\Delta_{\rho}^{p_{k}}(g-u_{\tau})\rangle_{\mathrm{L}^{2}(\mu)}$
@@
Since the choice of $T_{n}$ is arbitrary among all maps satisfying $(T_{n})_{\#}\mu=\mu_{n}$, we take the optimal one minimizing $\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{2}(\mu)}$. By the probabilistic transport bound of [27], this distance satisfies

### v2

@@ 3.4.1 Fully supervised problem with HOHL regularization
(5) $\|u_{n,\tau}^{(\mathbf{y}_{n})}-g|_{\Omega_{n}}\|_{\mathrm{L}^{2}(\mu_{n})}$ $\leq C\left[\sum_{k=1}^{q}\lambda_{k}\frac{\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\left(\varepsilon_{n}^{(k)}\right)^{2p_{k}}}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}+\frac{\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}+\tau\left(1+\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}\right)\right]$
where $u_{n,\tau}^{(\mathbf{y}_{n})}$ is the minimizer of $\mathcal{R}_{n,\tau}^{(\mathbf{y}_{n})}$.
We note from Propositions 4.3 and 4.4 that the variance term scales as
$\sum_{k=1}^{q}\lambda_{k}\frac{(\varepsilon_{n}^{(1)})^{2p_{1}}}{(\varepsilon_{n}^{(k)})^{2p_{k}}}\Big(\frac{\log(n)}{n(\varepsilon_{n}^{(k)})^{d}}\Big)^{1/2}+\frac{(\varepsilon_{n}^{(1)})^{2p_{1}}}{\tau}+\tau\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)},$
while, by the proof of Theorem 3.1, the bias term is of order $\tau$. This result highlights the nontrivial interplay between the multiple length-scales $\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$ and the regularization parameter $\tau$, which jointly determine the convergence behavior of the discrete minimizers. Understanding this relationship provides practical guidance for choosing these parameters to achieve an appropriate balance between bias, variance, and computational cost. Finally, the theorem generalizes known rates for graph-based learning: when $q=1$, we recover the convergence rates established in [trillos2022rates, Corollary 1.8], thereby situating our result within and extending the existing theoretical framework on graphs.
Remark 3.2 (Optimal regularization parameter $\tau$).
For fixed bandwidth parameters $\varepsilon_{n}^{(k)}$, the last two terms in (5) have the form $A/\tau+B\tau$, where $A=(\varepsilon_{n}^{(1)})^{2p_{1}}\text{ and }B=1+\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}.$ Minimizing this expression over $\tau>0$ yields the optimal choice $\tau^{\star}=\frac{(\varepsilon_{n}^{(1)})^{p_{1}}}{\sqrt{B}}$ and, substituting the latter in (5) gives
$\|u_{n,\tau^{\star}}^{(\mathbf{y}_{n})}-g|_{\Omega_{n}}\|_{L^{2}(\mu_{n})}\leq C\left[\sum_{k=1}^{q}\lambda_{k}\frac{\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\left(\varepsilon_{n}^{(k)}\right)^{2p_{k}}}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}+2\left(\varepsilon_{n}^{(1)}\right)^{p_{1}}\sqrt{1+\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}}\right].$
@@ 3.4.2 Truncated energies
Next, we show that we can use the truncated version of HOHL in practice. In fact, going beyond heuristics, the below results shows that truncated energies converges to the same continuum energy as the full energy (see [weihs2025Hypergraphs, Theorem 3.5]) This signifies that for large enough $n$, truncated and full energies will lead to arbitrarily close minimizers.
@@
Moreover, we note from Lemma 4.6 that the limiting continuum energy in Theorem 3.3
$\sum_{i=1}^{\infty}\Bigl(\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}\Bigr)\langle v,\psi_{i}\rangle_{\mathrm{L}^{2}(\mu)}^{2}$
is equal to $\langle v,\mathcal{L}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu)}$. In particular, this implies that $\mathcal{L}^{(q)}$ could also be defined directly through its spectrum: it shares the same eigenfunctions as $\Delta_{\rho}$, while its eigenvalues are given by functions of those of $\Delta_{\rho}$. This places our approach firmly within the framework of spectral kernel learning, where it is common to regularize with operators derived from the Laplacian. Spectral learning—and its analysis through reproducing kernel Hilbert space techniques—has been shown to yield powerful results for uncertainty quantification, enabling explicit bounds on expected error as well as estimates of prediction variance in semi-supervised learning (see [zhang2005spectral] and references therein).
@@ 3.4.3 Non-geometrical setting
All of the preceding results focused on applying HOHL within the geometric setting. The following result extends the analysis to arbitrary hypergraphs, demonstrating that the matrix $\mathcal{L}^{(q)}_{\mathrm{dis}}$ can be interpreted as the Laplacian of a specially constructed graph (which may be signed [signedGraphs]).
@@
This result is particularly noteworthy as it implies that standard numerical techniques developed for Laplace learning are directly applicable to HOHL. These include spectral truncation (see Theorem 3.3), Nyström extensions [fowlkes2004spectral], conjugate gradient methods for Laplacian inversion, and more. Moreover, it suggests that HOHL can function as a drop-in replacement for Laplace learning within existing machine learning pipelines. To demonstrate this in practice, Section 5 presents active learning experiments where the HOHL matrix $\mathcal{L}^{(q)}_{n}$ defines a Gaussian prior over functions.
We can extend the HOHL energy (2) to non-geometric datasets, where geometric embeddings for the vertices are unavailable and, for example, the weight models described in Section 3.2 do not apply. In such settings, the standard feature-based hypergraph construction, e.g. [scholkopfHyper2006, TVHg], forms a hyperedge among all nodes that share a common categorical feature value. Each hyperedge is also assigned unit weight.
@@
Continuum limit of HOHL $\mathrm{W}^{p_{q},2}$ semi-norm [weihs2025Hypergraphs] No natural continuum limit
Characterization of well/ill-posedness of HOHL in SSL ✓([weihs2025Hypergraphs, Theorem 3.5])
@@ 4.1 Fully supervised problem with HOHL regularization
First, the aim is to show the analogue of [trillos2022rates, Proposition 2.1] and to this purpose, we define
@@
Let $E=\cap_{r=1}^{q}E_{r}$ be the set of events such that the above inequalities hold for all $\leq r\leq q$. By [trillos2022rates, Lemma 2.2], we know that $\mathbb{P}(E_{r})\geq 1-2ne^{-c(r)n\left(\varepsilon_{n}^{(r)}\right)^{d}}$. Hence,
@@
where we used the fact that $\left(W_{n,\varepsilon_{n}^{(k)}}\right)_{ij}\leq C\left(\varepsilon_{n}^{(k)}\right)^{-d}$ and (8) for (10) and the fact that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ for (11). This implies that $n^{\ell-1}(\varepsilon_{n}^{(1)})]^{-1}q_{j}^{i}$ are sub-Gaussian and satisfy the same inequalities in the Birnbaum-Orlicz norm as in [trillos2022rates, Lemma 2.4]. By applying the same Hoeffding inequality as in the latter, for any $t>0$, we therefore obtain
@@
The term $T_{-r_{\chi}}$ contains $p_{k}-r_{\chi}$ factors, each equal to either $D_{n,\varepsilon_{n}^{(k)}}$ or $W_{n,\varepsilon_{n}^{(k)}}$. Using the operator-norm bounds from [trillos2022rates, Lemma 2.3], we have $\|T_{-r_{\chi}}\|_{\mathrm{op}}\leq(Cn)^{p_{k}-r_{\chi}}$. This implies that
@@
$\qquad\qquad=\left\|\left[\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\prod_{i=1}^{p_{k}}D_{n,\varepsilon_{n}^{(k)}}^{\chi_{i}}(-W_{n,\varepsilon_{n}^{(k)}})^{1-\chi_{i}}\right]\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
$\qquad\qquad\leq\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\left\|\prod_{i=1}^{p_{k}}D_{n,\varepsilon_{n}^{(k)}}^{\chi_{i}}(-W_{n,\varepsilon_{n}^{(k)}})^{1-\chi_{i}}\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
$\qquad\qquad\leq\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}\|T_{-r_{\chi}}\|_{\mathrm{op}}\left\|W_{n,\varepsilon_{n}^{(k)}}D_{n,\varepsilon_{n}^{(k)}}^{r_{\chi}-1}\tilde{w}_{n}\right\|_{\mathrm{L}^{2}(\mu_{n})}$
(20) $\qquad\qquad\leq C\sum_{\begin{subarray}{c}\chi\in\{0,1\}^{p_{k}}\\ \chi\neq(1,\dots,1)\end{subarray}}n^{p_{k}-r_{\chi}}\frac{n^{r_{\chi}}\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}$
(21) $\qquad\qquad=C(2^{p_{k}}-1)\frac{n^{p_{k}}\left(\varepsilon_{n}^{(1)}\right)^{2p_{1}}}{\tau}\left(\frac{\log(n)}{n\left(\varepsilon_{n}^{(k)}\right)^{d}}\right)^{1/2}$
@@
where we used the fact that there exists $C_{1}\leq n^{-1}d_{n,i,\varepsilon_{n}^{(r)}}\leq C_{2}$ for all $1\leq i\leq n$ and $1\leq r\leq q$ for (23) and the fact that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ for (24). Let $A$ be the event such that (15) holds. Then, arguing as in [Trillos3, Lemma 2.6], we can show that $\mathbb{P}(A\,|\,E)\geq 1-Cn^{-\alpha}$. Analogously to the proof of Lemma 4.1, we conclude that $\mathbb{P}(A)\geq 1-Cn^{-\alpha}$. ∎
@@
$\left\langle\frac{1}{2}\nabla\mathcal{R}_{n,\tau}^{(\mathbf{a}_{n})}(v^{(1)}_{n})-\frac{1}{2}\nabla\mathcal{R}_{n,\tau}^{(\mathbf{a}_{n})}(v_{n}^{(2)}),v_{n}^{(1)}-v_{n}^{(2)}\right\rangle_{\mathrm{L}^{2}(\mu_{n})}=\|v_{n}^{(1)}-v_{n}^{(2)}\|_{\mathrm{L}^{2}(\mu_{n})}^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\tau\sum_{k=1}^{q}\lambda_{k}\left\langle\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}\left(v_{n}^{(1)}-v_{n}^{(2)}\right),v_{n}^{(1)}-v_{n}^{(2)}\right\rangle_{\mathrm{L}^{2}(\mu_{n})}.$
@@
Furthermore, by first order optimality and (16), we have that
@@
where $u_{n,\tau}^{(\mathbf{g}_{n})}$ and $u_{\tau}$ are the minimizers of $\mathcal{R}_{n,\tau}^{(\mathbf{g}_{n})}$ and $\mathcal{R}_{\infty,\tau}^{(g)}$ respectively.
@@
then $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq\|\mathbf{a}_{n}\|_{\mathrm{L}^{2}(\mu_{n})}$. Indeed, by Proposition 3.4, we know that $\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}$ is a graph Laplacian, so we can apply the same proof as in [trillos2022rates, Lemma 2.14] with the eigenpairs of $\mathcal{L}_{n}^{(q)}=\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}$ to deduce (31).
@@
Let $E_{k}$ be the event such that [trillos2022rates, Theorem 2.8] holds for $\varepsilon_{n}^{(k)}$: we have
@@
where we used the fact that $u_{\tau}|_{\Omega_{n}}-u_{n,\tau}^{(\mathbf{g}_{n})}$ satisfies (31) with $\mathbf{a}_{n}=\tau\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}^{(k)}}^{p_{k}}-\sum_{k=1}^{q}\lambda_{k}\Delta_{\rho}^{p_{k}}\right)u_{\tau}|_{\Omega_{n}}$ for (36) and [trillos2022rates, Theorem 2.8] for (37).
@@
where we used (39) for (40). By [Stuart, Lemma 17], there exists $c$ and $C$ such that $c\|h\|_{\mathrm{W}^{s,2}(\Omega)}\leq\|h\|_{\mathcal{H}^{s}(\Omega)}\leq C\|h\|_{\mathrm{W}^{s,2}(\Omega)}$ for all $h\in\mathcal{H}^{s}(\Omega)$. From the above, we therefore deduce that $\|u_{\tau}\|_{\mathrm{W}^{s,2}(\Omega)}\leq C\|g\|_{\mathrm{W}^{s,2}(\Omega)}$. Finally, by Morrey’s inequality [leoni2017first], for $s$ sufficiently large there exists C>0C’>0 such that
@@
(43) $\qquad\qquad=\langle\nabla\mathcal{R}_{\infty}^{(g)}(g),g-u_{\tau}\rangle_{\mathrm{L}^{2}(\mu)}-\|g-u_{\tau}\|_{\mathrm{L}^{2}(\mu)}^{2}-\tau\sum_{k=1}^{q}\lambda_{k}\langle g-u_{\tau},\Delta_{\rho}^{p_{k}}(g-u_{\tau})\rangle_{\mathrm{L}^{2}(\mu)}$
@@
Since the choice of $T_{n}$ is arbitrary among all maps satisfying $(T_{n})_{\#}\mu=\mu_{n}$, we take the optimal one minimizing $\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{2}(\mu)}$. By the probabilistic transport bound of [fournier2015], this distance satisfies